OpenStudy (anonymous):

how do i find the inverse of C(n) = 28n+5700

OpenStudy (cwrw238):

make n in the subject of the function replace C(n) by n and n by C(-1) n

OpenStudy (anonymous):

n = 28(c(-1)) + 5700

OpenStudy (cwrw238):

you've done the replacement before making n the subject making n the subject means rearranging the formula to get n =

OpenStudy (cwrw238):

C(n) = 28n+5700 28n = C(n) - 5700 can you continue?

OpenStudy (anonymous):

ah. divide both sides by 28 . and ill have my inverse?

OpenStudy (cwrw238):

then you'll have n = whatever then you do the replacements

OpenStudy (anonymous):

i think i understand... but just to be clear.. am i just replacing n with C^-1(n)

OpenStudy (cwrw238):

thats part of it - also replace C(n) on the righthand side by n

OpenStudy (anonymous):

ok. thank you so much

OpenStudy (cwrw238):

yw